# Tackling GenAI Copyright Issues: Originality Estimation and Generecization

### v5

@@ 2 Performance of PREGen on COPYCAT
Simply rewriting the original prompt to a clean promt often fails to prevent generation of copyrighted characters [he2024fantasticcopyrightedbeastsnot], which indicates that the clean prompt still contains keywords and phrases that are strongly associated with these characters (see Table 8 for some examples of clean prompts). The additional prompt rewriting step using tamplate $\tau$ in PREGen introduces diversity in prompts to mitigate this issue.
@@
“Judy Hopps” “Create an image of an anthropomorphic rabbit character. She should be tall and slim, with a light grey fur and big, bright blue eyes. She should be wearing a dark blue police uniform with a badge, black boots, and a utility belt. She should be standing in a confident pose, showcasing her determination and ambition.”
“Mario” “Create an image of an animated plumber character. He should have a robust physique, a thick black moustache, and wear a red hat with a white “M” on it. His outfit should include blue overalls over a red shirt. He should have a joyful expression and be holding a green plumbing pipe.”
@@
Table 8: Sample clean prompts. Clean prompts obtained by rewriting direct anchors (character names) using the template system prompt $t$ are shown. Clean prompts appear to often contain keywords and phrases that trigger generation of copyrighted characters.

### v6

@@ 2 Performance of PREGen on COPYCAT
Simply rewriting the original prompt to a clean prompt often fails to prevent generation of copyrighted characters [he2024fantasticcopyrightedbeastsnot], which indicates that the clean prompt still contains keywords and phrases that are strongly associated with these characters (see Table 8 for some examples of clean prompts). The additional prompt rewriting step using template $\tau$ in PREGen introduces diversity in prompts to mitigate this issue.
@@
“Judy Hopps” “Create an image of an anthropomorphic rabbit character. She should be tall and slim, with a light gray fur and big, bright blue eyes. She should be wearing a dark blue police uniform with a badge, black boots, and a utility belt. She should be standing in a confident pose, showcasing her determination and ambition.”
“Mario” “Create an image of an animated plumber character. He should have a robust physique, a thick black mustache, and wear a red hat with a white “M” on it. His outfit should include blue overalls over a red shirt. He should have a joyful expression and be holding a green plumbing pipe.”
@@
Table 8: Sample clean prompts. Clean prompts obtained by rewriting direct anchors (character names) using the template system prompt $t$ appear to often contain keywords and phrases that trigger generation of copyrighted characters.